..17n-15+n=2n-8-0.4n can u please show me how to check it please and thanks
Answer by checkley71(8403)

A MUCH EASIER PROBLEM TO SOLVE AS YOU CAN SEE!!!!
.7N-15+N=2N-8-.4N
.7N+N-2N+.4N=15-8
.1N=7
N=7/.1
N=70 ANSWER.
PROOF
.7*70-15+70=2*70-8-.4*70
49-15+70=140-8-28
104=104